# Enabling Large-Scale and High-Precision Fluid Simulations on Near-Term Quantum Computers

### v2

@@ Abstract
Quantum computational fluid dynamics (QCFD) offers a promising alternative to classical computational fluid dynamics (CFD) by leveraging quantum algorithms for higher efficiency. This paper introduces a comprehensive QCFD method implemented on a superconducting quantum computer, demonstrating successful simulations of steady Poiseuille flow and unsteady acoustic wave propagation. The Poiseuille flow simulation achieved a relative error of less than $0.2\%$, and the unsteady acoustic wave simulation solved a 5043-dimension matrix, marking the largest fluid simulation on a quantum computer to date. Our approach bridges quantum and classical computing, adapting to quantum hardware constraints and offering scalable solutions for large-scale CFD problems, which paves the way for practical applications of near-term quantum computers in computational science.
@@
Quantum computational fluid dynamics, Superconducting quantum computer, Variational quantum linear solver, Quantum state tomography, Noisy intermediate-scale quantum
@@
[5]organization = CAS Key Laboratory of Quantum Information, University of Science and Technology of China, addressline = Jinzhai Road 92, city = Hefei, postcode = 230026, country = China
@@ 1 Introduction
Computational fluid dynamics (CFD) has played a pivotal role in numerous engineering and scientific disciplines, providing invaluable insights into complex flow behavior. However, as computational demands increase and the complexity of fluid dynamics problems intensifies, the limitations of classical computing architectures become increasingly apparent. Quantum computational fluid dynamics (QCFD) emerges as a promising alternative, offering the potential to overcome these limitations and unlock unprecedented computational power [1]. At its core, QCFD employs quantum algorithms and quantum computing hardware to simulate fluid dynamics phenomena with higher efficiency. The specific methods of QCFD are diverse, including Lagrangian-type methods [2, 3], explicit Euler-type methods [4], implicit Euler-type methods [5, 6], and methods based on quantum neural networks [7]. Focusing on solving complicated fluid governing equations, the linear system of equations plays a fundamental role in most classical CFD methods. While solving linear systems is the most time-consuming part of these methods, quantum linear solvers (QLS) promise an exponential speedup since first proposed in Ref. [8]. Tab. 1 reports these developments on QLS.
Table 1: The history of the QLS of the direct method [9].
Reference Year Primary advancement Query complexity
Harrow [8] 2009 First quantum approach $\mathcal{O}\left(\kappa^{2}/\epsilon\right)$
Ambainis [10] 2010 Variable-time amplitude amplification $\mathcal{O}\left(\kappa\left(\log\kappa/\epsilon\right)^{3}\right)$
Childs [11] 2017 Fourier or Chebyshev fitting using LCU $\mathcal{O}\left(\kappa\mathrm{poly}\log\left(\kappa/\epsilon\right)\right)$
Subasi [12] 2019 Adiabatic randomization method $\mathcal{O}\left(\kappa\log\kappa/\epsilon\right)$
An [13] 2022 Time-optimal adiabatic method $\mathcal{O}\left(\kappa\mathrm{poly}\log\left(\kappa/\epsilon\right)\right)$
Lin [14] 2020 Zeno eigenstate filtering $\mathcal{O}\left(\kappa\log\left(\kappa/\epsilon\right)\right)$
Costa [9] 2022 Discrete adiabatic theorem $\mathcal{O}\left(\kappa\log\left(1/\epsilon\right)\right)$
Despite these prominent advancements, the implementation of QCFD still faces challenges based on near-term quantum devices, i.e., noisy intermediate-scale quantum (NISQ) devices [15]. These aforementioned QLS are designed in the context of the fault-tolerant quantum computation (FTQC) era [16, 17]. In a fault-tolerant quantum computer, redundant noisy physical qubits are employed to encode an ideal “logical qubit” without error [18]. However, only a few logical qubits can be demonstrated so far [19, 20], preventing most QLS from near-term practicability.
Some efforts are conducted to enable computing fluid flows on quantum computers. One methodology is to emulate the flow field as a quantum state, utilizing the quantum state evolution to emulate the flow simulation, as shown in [21, 22]. Another methodology tries a variational approach to solve the equation, such as employing variational quantum linear solvers (VQLS) [23, 24, 25] in CFD, including solving potential and Stokes flows, lid-driven cavity, and the advection-diffusion equation [26, 27, 28, 29]. However, current implementations are limited to small sizes due to the inherent limitations of near-term quantum computers, which are significantly affected by noise. Additionally, encoding the entire flow field into a quantum state and reading it out into classical data using full quantum tomography is highly inefficient. These noise and readout issues impede the scalability of existing methods to larger problems.
In this paper, we propose a comprehensive method that enables large-scale and high-precision fluid simulation on near-term quantum computers. We implement it on a superconducting quantum computer, successfully simulating steady Poiseuille flow and unsteady acoustic wave propagation to showcase the ability of our method. In the Poiseuille flow simulation, we achieve less than $0.2\%$ relative error for showcasing the accuracy of a single step. In the unsteady acoustic wave propagation simulation, we successfully simulated 1681 grid points. Each time step solves a 5043-dimension matrix, which is the largest fluid simulation on the quantum computer reported so far.
To mitigate experimental error arising from noisy devices and approximation error, we introduce Iterative-QLS, by introducing iterations to enhance the precision rather than employing a single VQLS. This cycle aims to diminish residual error by achieving synthetic accuracy levels even lower than the physical accuracy. To overcome the problem size constraints imposed by intermediate-scale quantum processors, the subspace method is employed. This approach enables the hybrid utilization of quantum computing and classical computing, enabling simulation on a larger flow field. Furthermore, to enhance the efficiency of extracting classical information from quantum states, we have developed the sparse tomography technique. In summary, we present an integrated QCFD protocol capable of deriving high-precision solutions for problems of arbitrary size.
@@
We regard our algorithm as a bridge between quantum and classical computers. According to our findings, our method can adapt to the quantum computer’s capabilities, enabling collaboration between quantum and classical computers. Although the current solution uses only a limited number of qubits due to hardware constraints, advancements in quantum computing offer the possibility of solving larger subspaces by investing more quantum resources. Our approach is also a general-purpose quantum-enhanced linear solver method, applicable to other potential applications in computational science, including but not limited to finite element analysis, computer graphics, and signal processing. With successful, high-precision demonstrations on large-scale CFD problems, this paper marks a step towards scalable quantum-enhanced solutions for CFD and paves the way for the practical application of near-term quantum computers.
@@ 2.1 Basic concepts in quantum computing
The basic information unit in quantum information is the qubit, which is analogous to a classical bit but with the ability to exist in a superposition of states. Mathematically, a single qubit state can be represented as a vector in 2-dimension Hilbert space:
$\left|\psi\right\rangle=\alpha\left|0\right\rangle+\beta\left|1\right\rangle,$
@@
In quantum computing, we perform various operations on qubits, which store quantum information, to accomplish specific computational tasks. Quantum computing has different computational models, such as quantum annealing, one-way quantum computing, and quantum circuits, each of which can theoretically be converted into the others. Quantum circuit model is one of the most common models of quantum computing. The basic unit of a quantum circuit is the quantum logic gate, with each gate representing a unitary matrix. Common quantum logic gates include single-qubit gates and multi-qubit gates. Single-qubit gates include the X gate and the Hadamard gate, with their specific matrix forms as follows:
$X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}.$
@@
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix},\quad\text{CZ}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix}.$
Quantum logic gates can process $2^{n}$-dimensional quantum state with $O(n)$ or $O(1)$ complexity, achieving coherent superposition and interference between each dimension of the target quantum state. These features of quantum logic gates can be utilized to construct quantum algorithms that accelerate the solution of specific problems. Numerous quantum algorithms have been proposed to accelerate the solution of various problems, with applications spanning mathematics, physics, chemistry, biology, computer science, and more. Typical problems include integer factorization, search, Hamiltonian simulation, matrix inversion, classification, prediction, and others. For more information, refer to the Quantum Algorithm Zoo.
@@ 2.2 The technical path of QCFD using QLS
In CFD, one must solve linear or nonlinear partial differential equations, such as the Navier-Stokes equations. These equations can be represented as
$\partial{Q}/\partial{t}=\mathcal{F}(Q),$ (1)
where $\mathcal{F}(Q)$ represents linear or nonlinear operators. The first step is discretization, which involves breaking down the equations into a form that can be handled computationally. Both time and space discretization schemes are employed, which can be either explicit or implicit. While explicit methods have strict stability requirements limiting the time step, implicit methods are preferred for their ability to handle larger time steps and provide numerical stability, particularly for stiff differential equations. Therefore, we use implicit methods, which are formulated as
$\frac{Q^{t+1}-Q^{t}}{\Delta t}=R(Q^{t+1}),$ (2)
where $t$ represents the time step, and $R(Q^{t+1})$ is the residual.
Following time discretization, spatial discretization is performed on the equations. This can be done using finite difference, finite element, or finite volume methods. The discretized equation is expressed as
$\frac{\vec{Q}^{t+1}-\vec{Q}^{t}}{\Delta}=\vec{R}(\vec{Q}^{t+1}),$ (3)
where $\vec{Q}^{t+1}\in\mathbb{R}^{n}$, and $n$ represents the dimension of the discretized space. Generally, when $\mathcal{F}(Q)$ is nonlinear, this leads to a nonlinear system of equations for $\vec{Q}^{t+1}$, necessitating a linearization procedure.
@@
The final step involves solving these linear equations using QLS. Several QLS algorithms have been developed, see Tab. 1. The selection of an appropriate QLS depends on the specific CFD problem, considering the problem’s scale and the quality of the quantum chip. By executing the QLS in conjunction with classical computations, we complete the solution of the CFD problem, leveraging the power of quantum computing to enhance computational efficiency.
2.3 Challenges of implementing CFD simulation on NISQ devices
Currently, several QCFD algorithms have been proposed, but there are still challenges when solving specific CFD problems on NSIQ devices. Here, we list and analyze these challenges.
@@ 2.3 Challenges of implementing CFD simulation on NISQ devices
Thirdly, due to the “no-cloning theorem” [38], transferring quantum states to classical computers is challenging because the quantum state collapses after each measurement. Reconstructing the quantum state requires rerunning the entire computation for each subsequent measurement. This process, known as “quantum tomography,” is time-consuming and presents a significant obstacle to realizing practical quantum applications [38]. Improving the efficiency and precision of quantum tomography methods is crucial for achieving quantum advantage. $l_{2}$-norm tomography [39] reconstructs quantum states by minimizing the average error between the measured data and the model prediction but is sensitive to large outliers, which can distort the reconstruction if not properly managed. In contrast, $l_{\infty}$-norm tomography [40] provides strong guarantees on the worst-case error, though it demands quantum random access memory, which is an infeasible component in the near term. For noisy quantum devices with limited resources, efficient quantum state tomography is an urgent need.
@@ 3.1 Overview
When solving large-scale problems, the number of qubits required by QCFD algorithms may exceed the chip’s capacity. Therefore, a key issue for QCFD algorithms is how to solve large-scale CFD problems on NISQ devices with a limited number of qubits. In classical algorithms, subspace methods can transform the solution of high-dimensional linear systems into a series of lower-dimensional linear systems. We can apply subspace methods to QCFD algorithms to address scalability issues. However, subspace methods require high-precision solutions for each linear system, imposing a new challenge on QCFD algorithms, called high precision challenge.
@@
To address these challenges, we implement a comprehensive approach, as illustrated in Fig. 1, to ensure compatibility on NISQ devices and high-precision solutions. First, we introduce the Iterative-QLS approach for each small problem, enabling a high-precision quantum solver on noisy and limited-scale devices. The cycle is repeated until the residual of the approximated solution is suppressed below an expected precision to meet the desired convergence condition demanded by the divide-and-conquer subspace method. Second, the less-noisy while limited-scale solutions for different Krylov subspaces are synthesized to derive the large-scale solution. To mitigate the quantum-classical information conversion bottleneck, a sparse tomography technique is developed to extract the desired information from the readout more efficiently.
@@ 3.2 Iterative-QLS: Improving the precision of noisy quantum linear solver via iterative scheme
Figure 2: Flowchart of Iterative-QLS
To improve solution accuracy, an iterative method is utilized. In each loop, we first derive an approximate solution $\tilde{\boldsymbol{x}}_{i}$ by training VQLS. The training process of VQLS is similar to a standard VQLS process, where we exclude the Hadamard-Test and replace it with a loss function based on the results of tomography. B.3 shows how to evaluate the loss function and its gradient. The tomography process is the sparse tomography proposed in this paper, with details provided in Sec. 3.4.
Secondly, we evaluate the optimal magnitude based on the principle of $l_{2}$ norm. To obtain the optimal magnitude, we propose the “principle of minimum $l_{2}$ norm”. Assuming the noisy quantum state of the solution to the system of equations provided by VQLS is denoted as $|y\rangle$, the magnitude $L_{y}$ needs to be determined using certain optimal strategies. Here, $L_{y}$ should be the solution of the following optimization problem
@@
1:Linear equations $A\boldsymbol{x}=\boldsymbol{b}$. Convergence monitor $\epsilon_{c}$. Convergence monitor of loss value $\epsilon_{l}$.
@@
6:     Train VQLS process until satisfying convergence criteria $\epsilon_{c}$.
@@
9:     $\boldsymbol{y}\leftarrow\lVert\boldsymbol{y}\rVert_{2}\cdot\hat{\boldsymbol{y}}$

### v3

@@ Abstract
Quantum computational fluid dynamics (QCFD) offers a promising alternative to classical computational fluid dynamics (CFD) by leveraging quantum algorithms for higher efficiency. This paper introduces a comprehensive QCFD method, including an iterative method “Iterative-QLS” that suppresses error in quantum linear solver, and a subspace method to scale the solution to a larger size. We implement our method on a superconducting quantum computer, demonstrating successful simulations of steady Poiseuille flow and unsteady acoustic wave propagation. The Poiseuille flow simulation achieved a relative error of less than $0.2\%$, and the unsteady acoustic wave simulation solved a 5043-dimensional matrix. We emphasize the utilization of the quantum-classical hybrid approach in applications of near-term quantum computers. By adapting to quantum hardware constraints and offering scalable solutions for large-scale CFD problems, our method paves the way for practical applications of near-term quantum computers in computational science.
@@
Quantum computational fluid dynamics, Superconducting quantum computer, Variational quantum linear solver, Noisy intermediate-scale quantum
@@
[4]organization = Chinese Aeronautical Establishment, city = Beijing, postcode = 100012, country = China
[…]
[5]organization = Xi’an Aeronautics Computing Technique Research Institute, AVIC, city = Xi’an, postcode = 710000, country = China
[…]
@@
[7]organization = CAS Key Laboratory of Quantum Information, University of Science and Technology of China, city = Hefei, postcode = 230026, country = China
@@ 1 Introduction
Computational fluid dynamics (CFD) is crucial for flow mechanism research and industrial design. Advances in computing have continually introduced new methodologies and paradigms in scientific computation. Initially, CFD methods based on the finite volume and finite difference methods were developed for CPU-based high-performance computers. Over the past decade, GPU-based heterogeneous computing has gained prominence, enabling complex and computationally expensive methods like the spectral element method [1], high-order finite volume/difference methods [2, 3], and machine learning-based methods [4, 5]. However, simulating extreme flow parameters and large-scale models remains unaffordable with the most advanced classical supercomputers, which are hitting performance bottlenecks due to the atomic scale of transistors. This has motivated the exploration of new computing paradigms, with quantum computing being one of the most promising.
Significant progress in quantum computing has encouraged innovative methods in CFD [6]. Currently, there are two major approaches to performing fluid simulations on quantum computers. One methodology emulates the flow field by a quantum state, simulating the fluid dynamics via Hamiltonian simulation using the hydrodynamic Schrödinger equation, as shown in [7, 8, 9]. Alternatively, a broader approach replaces specific classical algorithms in CFD with their quantum counterparts to tackle sub-problems [10, 11]. Of particular interest are linear systems of equations, which are central to many classical CFD methods [12, 13, 14] and often the most time-consuming and resource-intensive components, thus constraining large-scale fluid simulations. Quantum linear solvers (QLS), as demonstrated in Refs. [15, 16, 17], provide an exponential speedup over classical methods, making them a potential solution for performing fluid simulations on quantum computers.
Concurrently, advancements in quantum hardware have demonstrated advantages in specific problems experimentally [18, 19, 20]. With current quantum computers now having several hundred qubits, near-term quantum computing applications for challenging flow problems are foreseeable. Nevertheless, quantum computational fluid dynamics (QCFD) still faces significant challenges when implemented on near-term quantum devices, specifically noisy intermediate-scale quantum (NISQ) devices [21]. A major hurdle stems from quantum errors. Most quantum algorithms are conceived within the context of the fault-tolerant quantum computation (FTQC) era [22, 23], wherein noisy physical qubits are redundantly encoded to form an error-free “logical qubit” [24]. However, only a few logical qubits have been successfully demonstrated to date [25, 26], implying that quantum errors will remain a significant concern for near-term quantum computers. Error suppression methods, such as quantum error mitigation (QEM) [27], are therefore crucial for practical quantum applications [28]. In the context of QCFD, the temporal evolution of flows means that computational errors can accumulate and amplify, potentially obscuring the true flow dynamics. Hence, ensuring computational accuracy is paramount.
Despite these challenges, several noteworthy contributions have been made to achieve fluid simulations on NISQ devices. One idea is to achieve a high-quality solution by utilizing high-fidelity qubits, demonstrated by two-dimensional unsteady flows discretized spatially with 1024 grid points performed on a well-calibrated superconducting quantum computer, achieving the largest quantum fluid simulation reported so far [8]. Another idea is to find error-resilient quantum methods, introducing the variational framework to solve linear systems on near-term quantum computers, such as variational quantum linear solvers (VQLS) [29, 30, 31]. This approach has been applied to address problems such as potential flows, Stokes flows, lid-driven cavities, and the advection-diffusion equation [32, 33, 34, 35, 36, 37]. These works successfully demonstrate the feasibility of QCFD on NISQ devices. However, current methods still face scalability issues on near-term quantum computers, where quantum errors still hinder the current method from computing a larger-scale fluid flow. Therefore, a more effective way to suppress the quantum error should be proposed to enable large-scale quantum fluid simulation.
@@
In this paper, we propose a comprehensive method that enables large-scale and high-precision fluid simulation on near-term quantum computers. To demonstrate the ability of the method, the steady Poiseuille flow and unsteady acoustic wave propagation are simulated on a superconducting quantum computer. This algorithm can be regarded as a bridge between quantum and classical computers. According to our findings, our method can adapt to the quantum computer’s capabilities, enabling collaboration between quantum and classical computers. Although the current solution uses only a limited number of qubits due to hardware constraints, advancements in quantum computing offer the possibility of investing more quantum resources to achieve quantum speedup. Our approach is also a general-purpose quantum-enhanced linear solver method, applicable to other potential applications in computational science, including but not limited to finite element analysis, computer graphics, and signal processing. With successful, high-precision demonstrations of large-scale CFD problems on a real quantum computer, this paper marks a step towards scalable quantum-enhanced solutions for CFD and paves the way for the practical application of near-term quantum computers.
The remaining part is organized as follows. In Sec. 2, basic concepts in quantum computing and the technical path of QCFD using QLS, as well as challenges of implementing fluid simulation on near-term devices, are introduced. Sec. 3 presents the details of our proposed method. Flow simulation experiments on a real quantum computer are shown in Sec. 4. Concluding remarks are given in the final section.
@@ 2.1 Basic concepts in quantum computing
The basic information unit in quantum information is the qubit, which is analogous to a classical bit but with the ability to exist in a superposition of states. Mathematically, a single qubit state can be represented as a vector in two-dimensional Hilbert space:
$\left|\psi\right\rangle=\alpha\left|0\right\rangle+\beta\left|1\right\rangle,$ (1)
@@
In quantum computing, we perform various operations on qubits, which store quantum information, to accomplish specific computational tasks. Quantum computing has different computational models, such as quantum annealing, one-way quantum computing, and quantum circuits, each of which can theoretically be converted into the others [38]. Quantum circuit model is one of the most common models of quantum computing. The basic unit of a quantum circuit is the quantum logic gate, with each gate representing a unitary matrix. Common quantum logic gates include single-qubit gates and multi-qubit gates. Single-qubit gates include the X gate and the Hadamard gate, with their specific matrix forms as follows:
$X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}.$ (2)
@@
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix},\quad\text{CZ}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix}.$ (3)
A quantum logic gate can process a $2^{n}$-dimensional quantum state with one step, achieving coherent superposition and interference between each dimension of the target quantum state. These features of quantum logic gates can be utilized to construct quantum algorithms that accelerate the solution of specific problems. Numerous quantum algorithms have been proposed to accelerate the solution of various problems, with applications spanning mathematics, physics, chemistry, biology, computer science, and more. Typical problems include integer factorization, search, Hamiltonian simulation, matrix inversion, classification, prediction, and others. For more information, refer to the Quantum Algorithm Zoo.
@@ 2.2 The technical path of QCFD using QLS
In CFD, one must solve linear or nonlinear partial differential equations, such as the Navier-Stokes equations. These equations can be represented as:
$\partial{\boldsymbol{Q}}/\partial{t}=\mathcal{F}(\boldsymbol{Q}),$ (4)
where $\boldsymbol{Q}$ is the variable vector, and $\mathcal{F}(\boldsymbol{Q})$ represents linear or nonlinear operators. The temporal and spatial derivatives are discretized with explicit or implicit schemes. While explicit methods have strict stability requirements limiting the time step, implicit methods are preferred for their ability to handle larger time steps and provide numerical stability, particularly for stiff differential equations. Therefore, we use implicit methods, which are formulated as:
$\frac{\boldsymbol{Q}^{t+1}-\boldsymbol{Q}^{t}}{\Delta t}=R(\boldsymbol{Q}^{t+1}),$ (5)
where $\Delta t$ represents the time step, and $R(\boldsymbol{Q}^{t+1})$ is the residual. Generally, when $\mathcal{F}(\boldsymbol{Q})$ is nonlinear, this leads to a nonlinear system of equations for $\boldsymbol{Q}^{t+1}$, necessitating a linearization procedure.
@@
The final step involves solving these linear equations using QLS. The selection of an appropriate QLS depends on the specific CFD problem, considering the problem’s scale and the quality of the quantum chip. By executing the QLS in conjunction with classical computations, we complete the solution of the CFD problem, leveraging the power of quantum computing to enhance computational efficiency.
2.3 Challenges of implementing fluid simulation on near-term devices
Currently, several QCFD algorithms have been proposed, but there are still challenges when solving specific CFD problems on near-term quantum computers, i.e. NISQ devices. Here, we list and analyze these challenges.
@@ 2.3 Challenges of implementing fluid simulation on near-term devices
Thirdly, due to the “no-cloning theorem” [38], transferring quantum states to classical computers is challenging because the quantum state collapses after each measurement. Extracting the quantum state to classical data requires rerunning the entire computation for each subsequent measurement. This process, known as “quantum tomography”, is time-consuming and presents a significant obstacle to realizing practical quantum applications [38]. Improving the efficiency and precision of quantum tomography methods is crucial for achieving quantum advantage. $l_{2}$-norm tomography [47] reconstructs quantum states by minimizing the average error between the measured data and the model prediction but is sensitive to large outliers, which can distort the reconstruction if not properly managed. In contrast, $l_{\infty}$-norm tomography [48] provides strong guarantees on the worst-case error, though it demands quantum random access memory, which is an infeasible component in the near term. For noisy quantum devices with limited resources, efficient quantum state tomography is an urgent need.
@@ 3.1 Overview
When one performs large-scale fluid simulations on NISQ devices, the number or quality of qubits required by quantum fluid simulations may exceed the quantum computer’s capability. Therefore, a key issue is how to simulate large-scale fluid on NISQ devices with a limited number of qubits. In classical algorithms, subspace methods can transform the solution of original linear systems into a series of lower-dimensional linear systems. Similarly, we can apply subspace methods to QLS to address scalability issues on a quantum computer. However, subspace methods require high-precision solutions for each linear system, without which the original linear system may fail to converge, imposing new challenges on QLS.
@@
To address these challenges, we implement a comprehensive approach to simulate a large-scale fluid flow with high precision on near-term quantum computers, as illustrated in Fig. 1. First, we introduce an iterative approach for each small linear problem, called Iterative-QLS, to enable a high-precision quantum solver on noisy and limited-scale devices. The cycle is repeated until the residual of the approximated solution is suppressed below an expected precision to meet the desired convergence condition demanded by the divide-and-conquer subspace method. Second, the less-noisy while limited-scale solutions for different Krylov subspaces are synthesized to derive the large-scale solution. To mitigate the quantum-classical information conversion bottleneck, a sparse tomography technique is developed to extract the desired information from the readout more efficiently.
@@ 3.2 Iterative-QLS: Improving the precision of noisy quantum linear solver via iterative scheme
Figure 2: Flowchart of Iterative-QLS.
To fulfill the requirements of performing large-scale flow simulation, improving the precision of a single VQLS becomes the first step. Here, we propose Iterative-QLS by introducing iterations beyond each VQLS, as shown in Fig. 2. In each loop, we first derive an approximate solution $\tilde{\boldsymbol{x}}_{i}$ by executing a VQLS. We employ a variant of VQLS which is similar to a standard VQLS process, where we exclude the Hadamard-Test and replace it with a loss function based on the results of tomography. B.3 shows how to evaluate the loss function and its gradient.
After executing VQLS, the tomography process should be executed to extract the data from the quantum state. In this paper, we propose an efficient method to extract a real-valued quantum state, named “Sparse tomography” which is introduced in the subsequent section 3.4.
Secondly, we evaluate the optimal magnitude based on the principle of $l_{2}$ norm. To obtain the optimal magnitude, we propose the “principle of minimum $l_{2}$ norm”. Assuming the noisy quantum state of the solution to the system of equations provided by VQLS is denoted as $|y\rangle$, the magnitude $L_{y}$ needs to be determined using certain optimal strategies. Here, $L_{y}$ should be the solution of the following optimization problem:
@@
1:Linear equations $A\boldsymbol{x}=\boldsymbol{b}$. Convergence criteria $\epsilon_{c}$. Convergence criteria of loss value $\epsilon_{l}$.
@@
6:     Train VQLS process until satisfying the precision criteria $\epsilon_{l}$.
@@
9:     $\boldsymbol{y}\leftarrow L_{y}\cdot\hat{\boldsymbol{y}}$